# Variational Learning Finds Flatter Solutions at the Edge of Stability

### v1

@@ Abstract
Variational Learning (VL) has recently gained popularity for training deep neural networks and is competitive to standard learning methods. Part of its empirical success can be explained by theories such as PAC-Bayes bounds, minimum description length and marginal likelihood, but there are few tools to unravel the implicit regularization in play. Here, we analyze the implicit regularization of VL through the Edge of Stability (EoS) framework. EoS has previously been used to show that gradient descent can find flat solutions and we extend this result to VL to show that it can find even flatter solutions. This is obtained by controlling the posterior covariance and the number of Monte Carlo samples from the posterior. These results are derived in a similar fashion as the standard EoS literature for deep learning, by first deriving a result for a quadratic problem and then extending it to deep neural networks. We empirically validate these findings on a wide variety of large networks, such as ResNet and ViT, to find that the theoretical results closely match the empirical ones. Ours is the first work to analyze the EoS dynamics in VL.
@@ 1 Introduction
Despite these successes, the theoretical mechanisms behind the good performance of VL remain poorly understood. It is often assumed that simplistic Gaussian posteriors, such as those used currently for deep learning, may not be enough because they are poor approximations of the true posterior; lack of a good prior is another issue. Despite these concerns, VL shows good performance in practice. Theories such as minimum description length (Hinton and Van Camp, 1993; Hochreiter and Schmidhuber, 1997; Blier and Ollivier, 2018), PAC-Bayes (Dziugaite and Roy, 2017; Zhou et al., 2019; Lotfi et al., 2024) and marginal likelihood (Smith and Le, 2017; Immer et al., 2021) can partially explain the success but these theories say little about the regularizing properties of the learning algorithm. Similarly to deep learning, the presence of implicit regularization is likely also at play in VL, but few tools exist (if any) to unravel these effects.
@@ Appendix A Related Works
Injecting noise into the parameters within gradient descent has several desirable features such as escaping saddle points Jin et al. [2017], Reddi et al. [2018] and local minima Zhu et al. [2019], Nguyen et al. [2019]. In fact it has been widely observed empirically that the SGD noise has an important role to play to find flat minima. Noise with larger scale Zhu et al. [2019], Nguyen et al. [2019], Zhang et al. [2019], Smith et al. [2020], Wei and Schwab [2019] and heavy-tail Simsekli et al. [2019], Panigrahi et al. [2019], Nguyen et al. [2019], Wang et al. [2022a] drives the optimization trajectories towards flatter minima. Orvieto et al. [2023] demonstrated through stochastic Taylor expansion that injecting Gaussian noise into parameters before a gradient step implicitly regularizes by penalizing the curvature of the loss, while Orvieto et al. [2022] showed that anticorrelated noise in Perturbed Gradient Descent (PGD) specifically penalizes the trace of the Hessian. Zhang et al. [2024a] showed that injecting noise in opposite directions to the weight space leads to a better regularization of the trace of the Hessian.
@@
Cohen et al. [2021], building on the work of Jastrzebski et al. [2020] empirically showed that for full batch Gradient Descent (GD) with constant step-size ($\rho$), the operator norm of the Hessian (also termed as sharpness) settles in a neighborhood of $2/\rho$. This threshold is termed as edge of stability (EoS) because gradient descent on a quadratic only converges if the sharpness is below $2/\rho$. Strikingly in complex neural network landscapes, sharpness settling around the stability limit $2/\rho$ (instead of diverging) indicates that presence of a self-stabilization mechanism Damian et al. [2023] (which is absent in a quadratic) that regularizes the sharpness near $2/\rho$. The common occurrence of this phenomenon across various tasks, architectures and initialization has inspired substantial research on edge of stability. For example EoS has been studied across several non-convex optimization problems Wang et al. [2022b, 2023a], Zhu et al. [2023], Chen and Bruna [2023], Agarwala et al. [2023], Arora et al. [2022], Lyu et al. [2022], Ahn et al. [2024], Wu et al. [2024], Ghosh et al. [2025], Even et al. [2024], Chen et al. [2024], Kreisler et al. [2023], Kalra et al. [2025], Zhu et al. [2024] and across other descent based optimizers such as SGD Lee and Jang [2023], Andreyev and Beneventano [2024], momentum Phunyaphibarn et al. [2024], sharpness aware minimization (SAM) Agarwala and Dauphin [2023], Long and Bartlett [2024] and Adam/RMSProp Cohen et al. [2022, 2025].
@@ Appendix B Proof of Theorem-1
We write the perturbed gradient as $\hat{\mathbf{g}}=\mathbf{g}+\boldsymbol{\xi}$, where $\boldsymbol{\xi}\sim\mathcal{N}\left(\mathbf{0},\frac{1}{N_{s}}\mathbf{Q}\boldsymbol{\Sigma}\mathbf{Q}\right)$. This expresses the update as the sum of a deterministic component $\mathbf{g}$ and a random fluctuation $\boldsymbol{\xi}$.
@@
where $a_{i}:=(\mathbf{v}_{i}^{\top}\mathbf{g})^{2}>0$. This is a cubic polynomial in $\lambda_{i}$ of the form

### v2

@@ Abstract
Variational Learning (VL) has recently gained popularity for training deep neural networks. Part of its empirical success can be explained by theories such as PAC-Bayes bounds, minimum description length and marginal likelihood, but there are few tools to unravel the implicit regularization in play. Here, we analyze the implicit regularization of VL through the Edge of Stability (EoS) framework. EoS has previously been used to show that gradient descent can find flat solutions and we extend this result to VL to show that it can find even flatter solutions. This is obtained by controlling the posterior covariance and the number of Monte Carlo samples from the posterior. These results are derived in a similar fashion as the standard EoS literature for deep learning, by first deriving a result for a quadratic problem and then extending it to deep neural networks. We empirically validate these findings on a wide variety of large networks, such as ResNet and ViT, to find that the theoretical results closely match the empirical ones. Ours is the first work to analyze the EoS dynamics in VL.
@@ 1 Introduction
Despite these successes, the theoretical mechanisms behind the good performance of VL remain poorly understood. It is often assumed that simplistic Gaussian posteriors, such as those used currently for deep learning, may not be enough because they are poor approximations of the true posterior; lack of a good prior is another issue. Despite these concerns, VL shows good performance in practice. Theories such as minimum description length (Hinton and Van Camp, 1993; Hochreiter and Schmidhuber, 1997; Blier and Ollivier, 2018), PAC-Bayes (Dziugaite and Roy, 2017; Zhou et al., 2019; Lotfi et al., 2024) and marginal likelihood (Smith and Le, 2017; Immer et al., 2021) can partially explain the success. In particular, PAC-Bayes theory provides a natural explanation for why flatter solutions may generalize better, see for instance the discussion in (Alquier et al., 2024, Section 3.3). However these theories say little about the regularizing properties of the learning algorithm. Similarly to deep learning, the presence of implicit regularization is likely also at play in VL, but few tools exist (if any) to unravel these effects.
@@ Appendix A Related Works
Injecting noise into the parameters within gradient descent has several desirable features such as escaping saddle points [Jin et al., 2017, Reddi et al., 2018] and local minima Zhu et al. [2019], Nguyen et al. [2019]. In fact it has been widely observed empirically that the SGD noise has an important role to play to find flat minima. Noise with larger scale [Zhu et al., 2019, Nguyen et al., 2019, Zhang et al., 2019, Smith et al., 2020, Wei and Schwab, 2019] and heavy-tail [Simsekli et al., 2019, Panigrahi et al., 2019, Nguyen et al., 2019, Wang et al., 2022a] drives the optimization trajectories towards flatter minima. Orvieto et al. [2023] demonstrated through stochastic Taylor expansion that injecting Gaussian noise into parameters before a gradient step implicitly regularizes by penalizing the curvature of the loss, while Orvieto et al. [2022] showed that anticorrelated noise in Perturbed Gradient Descent (PGD) specifically penalizes the trace of the Hessian. Zhang et al. [2024a] showed that injecting noise in opposite directions to the weight space leads to a better regularization of the trace of the Hessian.
@@
Cohen et al. [2021], building on the work of Jastrzebski et al. [2020] empirically showed that for full batch Gradient Descent (GD) with constant step-size ($\rho$), the operator norm of the Hessian (also termed as sharpness) settles in a neighborhood of $2/\rho$. This threshold is termed as edge of stability (EoS) because gradient descent on a quadratic only converges if the sharpness is below $2/\rho$. Strikingly in complex neural network landscapes, sharpness settling around the stability limit $2/\rho$ (instead of diverging) indicates that presence of a self-stabilization mechanism Damian et al. [2023] (which is absent in a quadratic) that regularizes the sharpness near $2/\rho$. The common occurrence of this phenomenon across various tasks, architectures and initialization has inspired substantial research on edge of stability. For example EoS has been studied across several non-convex optimization problems [Wang et al., 2022b, 2023a, Zhu et al., 2023, Chen and Bruna, 2023, Agarwala et al., 2023, Arora et al., 2022, Lyu et al., 2022, Ahn et al., 2024, Wu et al., 2024, Ghosh et al., 2025, Even et al., 2024, Chen et al., 2024, Kreisler et al., 2023, Kalra et al., 2025, Zhu et al., 2024] and across other descent based optimizers such as SGD [Lee and Jang, 2023, Andreyev and Beneventano, 2024], momentum [Phunyaphibarn et al., 2024], sharpness aware minimization (SAM) Agarwala and Dauphin [2023], Long and Bartlett [2024] and Adam/RMSProp [Cohen et al., 2022, 2025].
@@ Appendix B Proof of Theorem-1
We write the perturbed gradient as $\hat{\mathbf{g}}=\mathbf{g}+\boldsymbol{\xi}$, where $\boldsymbol{\xi}\sim\mathcal{N}\left(\mathbf{0},\frac{1}{N_{s}}\mathbf{Q}\boldsymbol{\Sigma}\mathbf{Q}\right)$. Here $\mathbf{g}$ denotes the gradient evaluated at the posterior mean, that is, $\nabla\ell(\boldsymbol{m}_{t})$. This expresses the update as the sum of a deterministic component $\mathbf{g}$ and a random fluctuation $\boldsymbol{\xi}$.
@@
where $a_{i}:=(\mathbf{v}_{i}^{\top}\mathbf{g})^{2}>0$  [^1] For GD, if $\mathbf{v}_{i}^{T}\mathbf{m}_{0}=0$, $a_{i}=0$ always holds. However, stochasticity ensures that $a_{i}\neq 0$ even when an $\mathbf{m}_{0}$ is chosen such that it is orthogonal to some $\mathbf{v}_{i}$’s.. This is a cubic polynomial in $\lambda_{i}$ of the form